OpenStudy (anonymous):

2[2/(2-x)][-2/(2-x)^2] my teacher wants us to simplify using only positive exponents but i don't even know where to start.

OpenStudy (amistre64):

start by expanding if need be; but then i cant see what a simplified form would mean

OpenStudy (anonymous):

and right now, all the exponents *are* positive...

OpenStudy (anonymous):

I don't see any negative exponents..

OpenStudy (amistre64):

\[\frac{2}{1}\frac{2}{(2-x)}\frac{-2}{(2-x)^2}\] might make it easier to read thru

OpenStudy (anonymous):

i think he means by the end all the exponents should be positive

OpenStudy (anonymous):

so...just multiply like you would any other fraction...

OpenStudy (amistre64):

-1*2*2*2 = -1*2^3 = -2^3 for the top right?

OpenStudy (anonymous):

2 x 2 x (-2) = -8 on top

OpenStudy (amistre64):

the bottom is just treating (2-x) is tho it was just some other number

OpenStudy (anonymous):

oh...i c. good call on the -2^3, amistre

OpenStudy (anonymous):

you're simplified form will be \[(\frac{?}{?})^{3}\]

OpenStudy (anonymous):

ok. that makes sense.. just multiply like a regular fraction and then try to cancel out and such, right?

OpenStudy (amistre64):

yep, but in this case nothing cancels

OpenStudy (amistre64):

it just morphs into one conglomerate mass :)